you have ten six-sided dice in a cup. If you shake the cup, toss all
of the dice on the table at one time, and count the number of threes, how many
threes would you expect to see?
What percentage of the time would you expect to see an outcome of five
threes?

Answer by stanbon(75887)

you have ten six-sided dice in a cup. If you shake the cup, toss all
of the dice on the table at one time, and count the number of threes, how many
threes would you expect to see?
---
Ans: (1/6)*10 = 10/6 = 5/3
----------------------------------
What percentage of the time would you expect to see an outcome of five
threes?
---
Binomial with n = 10 and p = 1/6
P(x=5) = 10C5(1/6)^5*(5/6)^5 = 0.013
=============
Cheers,
Stan H.
==========